OpenStudy (anonymous):

True or False? Nine-year-old Chico is 1/5 of his aunt Sofie's age. An equation to represent this problem is (1/5)a = 9. True False Nine-year-old Chico is 1/5 of his aunt Sofie's age. How old is Aunt Sofie? 45 50 65 55 Solve the following: You can get 12 color prints for $5.76. How much is this per print? $0.50 $0.49 $0.48 $0.47

OpenStudy (lgbasallote):

a/5 = 9 cross multiply.. a = 9(5)

OpenStudy (anonymous):

thequestion number 3 is .48$ right?

OpenStudy (lgbasallote):

5.76/12 = yes .48

OpenStudy (anonymous):

true and 45

OpenStudy (lgbasallote):

yes